# RISAR: RIS-assisted Human Activity Recognition with Commercial Wi-Fi Devices

### 27 Feb 2024 v1

@@ Abstract
Human activity recognition (HAR) is integral to smart homes, security, and healthcare. Existing systems face challenges such as the absence of line-of-sight links, insufficient details regarding the sensing subject, and inefficiencies in noise reduction and feature extraction from channel state information. To address these issues, this study builds a reconfigurable intelligent surface (RIS)-assisted passive human activity recognition (RISAR) system, compatible with commercial Wi-Fi devices. RISAR employs a RIS to augment the orientation and spatial diversity of Wi-Fi signals, thereby facilitating improved detection capabilities. A new denoising and feature extraction technique, the high-dimensional factor model, based on random matrix theory, is proposed for noise elimination and salient temporal feature extraction. On this basis, a dual-stream spatial-temporal attention network model is developed for assigning variable weights to different characteristics and sequences, mimicking human cognitive processes in prioritizing essential information. Experimental analysis shows that RISAR significantly outperforms existing HAR systems in accuracy and efficiency, achieving an average accuracy of 97.26%. These findings not only highlight the adaptability of the RISAR system but also underscore its potential as a robust solution for activity recognition across complex environments.
@@ I Introduction
Human activity recognition (HAR) is a crucial component in the Internet of Things (IoT) realm, particularly in smart homes and offices [1]. It entails the classification of human activities, thereby enabling applications in eldercare, healthcare, and home security [2, 3, 4, 5]. HAR methodologies can be systematically categorized into three distinct domains, predicated on the sensing modalities employed: vision-based HAR [6, 7, 8, 9], wearable device-based HAR [10, 2, 11, 12], and wireless sensing-based HAR [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25]. Vision-based HAR, which primarily employs high-resolution cameras, contends with impediments in non-line-of-sight (NLOS) scenarios and engenders privacy concerns [26]. Wearable device-based HAR, leveraging integrated sensors such as accelerometers, magnetometers, gyroscopes, barometers, and physiological sensors, faces the drawback of potential user inconvenience, particularly among the elderly and children, due to the necessity of continuous wear [26]. Wireless sensing-based HAR encompasses technologies like radar, radio frequency identification (RFID), and Wi-Fi signals. Radars, capable of detecting human motion without necessitating user-carried devices, offer resilience against variations in lighting conditions due to their broad bandwidth [27]. However, challenges in radar-based HAR systems include limitations in portability and elevated hardware costs. Conversely, RFID systems present a cost-effective alternative, requiring minimal hardware such as chip tags and antennas. Despite their economic advantages, RFID systems face robustness issues in intricate indoor environments, predominantly due to factors like multipath effect and temporal dynamics [28].
@@
Recent advancements in Wi-Fi signal-based HAR techniques have elicited significant interest in the field. These methodologies leverage channel state information (CSI) and multi-antenna configurations to effectively distinguish between various human activities, as demonstrated in systems such as WiSee [15], RT-Fall [16], CARM [17], WiAct [20], and CeHAR [24]. For instance, RT-Fall [16] utilized both amplitude and phase data from CSI measurements to identify falls, while CARM [17] applied CSI-speed and CSI-activity models for recognizing activities. Furthermore, WiAct [20] explored the correlation between body movement and CSI amplitude to classify activities. Despite the passive detection capabilities and the ease of deployment characterizing these systems, they face notable challenges, particularly in NLOS scenarios and in lack of spatial diversity. These limitations underscore the necessity for ongoing research and development in HAR technologies to achieve robust performance across a wider array of environmental conditions.
@@
In typical indoor environments, often comprised of multiple rooms linked to a single Wi-Fi access point (AP), Wi-Fi signals are frequently confronted with NLOS conditions. These scenarios are characterized by physical barriers and distance, which hinder both direct and reflected signal propagation paths between the transmitter and receiver. As a result, such environments experience elevated levels of noise, interference, and multipath effects [29]. Furthermore, traditional denoising and feature extraction methods, like principal component analysis (PCA), while effective for linear dimensionality reduction, may not adequately tackle the complexities associated with Wi-Fi signal data under these conditions [30]. Our empirical research indicates that PCA, in particular, delivers suboptimal performance in activity recognition within NLOS contexts. This finding underscores the imperative for investigating more sophisticated techniques to improve HAR performance in these demanding environments.
In this research, we present a novel device-free reconfigurable intelligent surface-assisted human activity recognition (RISAR) system that utilizes Wi-Fi signals, specifically designed for NLOS scenarios. The RISAR system comprises a pair of personal computers (PCs), with one functioning as the transmitter and the other as the receiver, in conjunction with RISs. Addressing the primary challenge, we integrate RIS technology into the HAR domain. Celebrated for its groundbreaking potential in the field of wireless communications, RIS enables the dynamic adaptation of propagation channels, thereby fostering an environment optimized for signal transmission [31]. This strategic enhancement facilitates the creation of multiple independent signal pathways, significantly boosting the accuracy of human activity recognition [32]. Moreover, the strategic deployment of RIS utilizes spatial diversity to effectively diminish the multipath effect, ensuring enhanced sensing performance [33, 34, 35]. To address the secondary challenge, we introduce a high-dimensional factor model for the extraction of activity features from the raw CSI measurements. Additionally, a specialized dual-stream spatial-temporal attention network (DS-STAN) architecture, expressly designed for the nuanced characteristics of CSI data, has been developed. This architecture is adept at processing and interpreting the complex patterns within CSI data, thereby enhancing our HAR system’s proficiency in the accurate identification and classification of human activities.
@@
    A RISAR system is built, which capitalizes on Wi-Fi signals, especially for NLOS conditions. The incorporation of RIS not only introduces unique signal pathways but also enhances spatial diversity, offering a significant improvement in signal quality and system performance. To the best of our knowledge, this is the first work to employ RIS technology for human activity recognition with commodity Wi-Fi devices.
  • $\bullet$
@@
    To enable the effective implementation of RISAR, a high-dimensional factor model approach is proposed for extracting activity features from CSI measurements. These extracted features are fed into the proposed DS-STAN model that integrates an attention mechanism to discriminate various human activities. This approach represents a sophisticated fusion of random matrix theory and deep learning techniques, offering enhanced accuracy in activity recognition.
  • $\bullet$
@@
    The RISAR system, operational with commercial Wi-Fi devices under both LOS and NLOS conditions, has undergone extensive evaluation. The empirical results highlight the system’s remarkable accuracy, reaching an average of 97.26% in LOS scenarios and 90.83% in NLOS configurations. Such performance not only illustrates RISAR’s robustness but also its adaptability across various environmental settings, signifying a substantial leap forward in HAR technology.
The remainder of this paper is as follows: Section II encompasses a comprehensive literature review of pertinent existing works. Section III presents the system information. Section IV details our experimental results, and Section V concludes the paper.
II Related Works
In this section, we summarize the related work on human activity recognition and Wi-Fi signal-based activity recognition.
II-A Human Activity Recognition
Current endeavors in HAR predominantly encompass three distinct categories: vision-based [6, 7, 8, 9], wearable device-based [10, 2, 11, 12], and wireless sensing-based systems [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25].
@@ II-A Human Activity Recognition
Vision-based methodologies primarily leverage cameras to capture image sequences and videos for activity recognition [6]. For instance, Kim et al. [7] introduced a depth video-based HAR system, utilizing skeletal joint features to discern the daily activities of the elderly in indoor settings. Additionally, Franco et al. [8] amalgamated skeleton and RGB data stream analyses to extract features for HAR. Despite their effectiveness, these systems are susceptible to lighting conditions and obstructions, and they raise significant security and privacy concerns.
@@
Wearable device-based HAR approaches utilize technologies such as smartphones [2], IoT devices [10], and sensors [11] for activity detection. Hassan et al. [2] developed a robust HAR system based on smartphone sensor data, capable of distinguishing transitional and non-transitional activities. Bhat et al. [10] employed a low-power IoT device for HAR, pioneering a framework for both online training and inference. Bianchi et al. [11] explored the integration of inertial measurement units (IMUs) with deep learning, where wearable sensors transmit data to a cloud service for activity monitoring. These methods, however, rely on active detection, necessitating the continuous wearing of the device.
In the realm of wireless sensing-based HAR, techniques utilize wireless signals for monitoring activities. Zhu et al. [13] proposed a distributed ultra-wideband (UWB) radar-based system, classifying motion events through spatio-temporal patterns of radar Doppler signatures. Oguntala et al. [14] introduced SmartWall, a system employing passive RFID tags for the recognition of sequential and concurrent activities. While these techniques obviate the need for wearing devices, they often require dedicated devices to be installed in the environment.
II-B Wi-Fi signal-based activity recognition
@@
In the evolving landscape of Wi-Fi signal-based technologies, a diverse array of systems have been developed for HAR. Yet, these systems frequently necessitate specialized devices and are often inadequate in NLOS scenarios. For instance, the WiSee system, as demonstrated by Pu et al. [15], confirmed the efficacy of Wi-Fi signals in detecting human motion in NLOS and through-the-wall environments, albeit relying on USRP-N210 devices rather than standard commercial Wi-Fi hardware. Yousefi et al. [18] employed long short-term memory (LSTM) networks to encode temporal information during feature learning, setting a new benchmark in HAR using CSI measurements. Similarly, Chen et al. [19] innovatively utilized raw CSI signals in conjunction with an attention-based bi-directional LSTM to ascertain action classes. Feng et al. [21] proposed a three-phase system for multiple human activity recognition, focusing on identifying the start and end points of activities in multi-subject scenarios. Xiao et al. [22] introduced a deep learning-based HAR approach that capitalizes on Wi-Fi signals, transforming segmentation tasks into classification challenges via a convolutional neural network (CNN) algorithm. Guo et al. [23] developed an encoder-decoder framework aimed at mitigating accuracy disparities among individuals in HAR applications. Lu et al. [24] in their CeHAR system, advocated for the amalgamation of CSI amplitude and phase features to enhance action detection efficiency. Lastly, Shalaby et al. [25] presented a deep learning model integrating CNNs, gated recurrent units (GRUs), and LSTMs, designed to extract high-dimensional and time sequence features.
@@
Despite their advantages in passive detection, ease of deployment, and cost-effectiveness, these systems generally do not consider HAR in NLOS environments. This is particularly pertinent in many indoor settings, where multiple rooms often connect to a single wireless signal AP, especially in residential environments. In such scenarios, issues like signal attenuation and multipath fading can significantly impair the accuracy of activity recognition.
III System Framework
The architecture of the RISAR system, as presented in Fig. 1, is comprised of several integral stages, namely: CSI data collection, data preprocessing, and human activity classification.
Figure 1: The framework of the proposed RISAR system.
III-A RISAR System
In contrast to traditional radio-frequency (RF) sensing methods, which are often constrained by their radio environments and limited in the number of transmission channels for multi-dimensional human activity information, our research introduces the RISAR system. This system is specifically designed for human activity recognition using RIS in NLOS environments. The data collection for activity recognition was conducted in three distinct environments, as depicted in Fig. 2. The RISAR system is comprised of two PCs, each equipped with an Intel Wi-Fi Link 5300 network interface card (NIC), functioning as the transmitter (Tx) and receiver (Rx). In environment A, the setup includes a PC with a horn antenna as the Tx and a PC with three antennas as the Rx. In environments B and C, both Tx and Rx are configured with three antennas each. When a human activity occurs within the transmission and reception range, the RISAR system is capable of automatically detecting it.
(a) Environment A.
(b) Floor plans of two experimental environments.
@@ III-A RISAR System
Under NLOS conditions, as illustrated in Fig.2LABEL:sub@F2-a, the experimental setup involves a Tx PC outfitted with an Intel 5300 NIC and a horn antenna, positioned 1 meter high, directing signals towards the RIS located at a corner 5 meters away, thereby establishing a LOS link. The RIS, placed at the same 1-meter height, is optimized using the DaS algorithm [36] to adjust its phase matrix, maximizing signal power at the Rx location. Concurrently, the Rx PC, also equipped with an Intel 5300 NIC and an omnidirectional antenna, is stationed on a 1-meter-high tripod. The efficacy of the RIS, as detailed in a prior study by the authors [37], is validated, with the designed RIS operating at 5.8 GHz yielding an improvement of over 10 dBi. For LOS conditions, the experimental setup, as delineated in Fig.2LABEL:sub@F2-b, incorporates data collection in a meeting room within a $12\times 8.7\,\text{m}^{2}$ area and an office room covering an $8\times 8.7\,\text{m}^{2}$ space. The Intel 5300 NIC in our experiments was configured to operate in the 5 GHz band, utilizing the 161-st channel with a 20 MHz channel bandwidth. The RIS board consists of a uniform planar array of $10\times 16$ units, totaling 160, and operates at a central frequency of 5.8 GHz.
@@ III-B Data Collection
CSI measurements are obtained from packets captured by the receiver. As delineated by Halperin et al. [38], each packet encompasses a CSI matrix of dimensions $N_{\text{Tx}}\times N_{\text{Rx}}\times N_{\text{sc}}$, where $N_{\text{Tx}}$ and $N_{\text{Rx}}$ signify the count of transmitting and receiving antennas, respectively, and $N_{\text{sc}}$ indicates the orthogonal frequency-division multiplexing (OFDM) subcarriers. Specifically, the Intel Wi-Fi Link 5300 NIC, which uses 56 subcarriers, exports CSI data for 30 subcarriers across three antennas, resulting in each packet incorporating $1\times 3\times 30$ CSI measurements for NLOS and $3\times 3\times 30$ for LOS scenarios.
In the proposed RISAR system, the CSI sampling rate is set at 1000 packets per second over a 5-second interval. We curate diverse datasets in environments labeled as RIS-corner (Environment A), Meeting room (Environment B), and Office (Environment C) to appraise the system’s capability in discerning a variety of daily activities. Each activity is executed 100 times by participants to amass a comprehensive dataset for analysis, as cataloged in Table I.
@@
RIS-corner Meeting room Office

### 20 Mar 2024 v3

@@ Abstract
Human activity recognition (HAR) holds significant importance in smart homes, security, and healthcare. Existing systems face limitations because of the insufficient spatial diversity provided by a limited number of antennas. Furthermore, inefficiencies in noise reduction and feature extraction from sensing data pose challenges to recognition performance. This study presents a reconfigurable intelligent surface (RIS)-assisted passive human activity recognition (RISAR) method, compatible with commercial Wi-Fi devices. RISAR leverages a RIS to enhance the spatial diversity of Wi-Fi signals, effectively capturing a wider range of information distributed across the spatial domain. A novel high-dimensional factor model based on random matrix theory is proposed to address noise reduction and feature extraction in the temporal domain. A dual-stream spatial-temporal attention network model is developed to assign variable weights to different characteristics and sequences, mimicking human cognitive processes in prioritizing essential information. Experimental analysis shows that RISAR significantly outperforms existing HAR methods in accuracy and efficiency, achieving an average accuracy of 97.26%. These findings underscore RISAR’s adaptability and potential as a robust activity recognition solution in real environments.
@@ I Introduction
Human activity recognition (HAR) is a crucial component in the Internet of Things (IoT) realm. HAR involves the classification of human activities, enabling applications in eldercare, healthcare, and home security [1, 2, 3, 4]. Recent advancements in Wi-Fi signal-based HAR techniques have highlighted their promising potential. These methodologies often leverage channel state information (CSI) to distinguish various human activities, as demonstrated in systems such as WiSee [5], RT-Fall [6], CARM [7], WiAct [8], and CeHAR [9]. Notably, RT-Fall [6] utilized both amplitude and phase data from CSI measurements to identify falls, while CARM [7] applied CSI-speed and CSI-activity models for recognizing activities. Furthermore, WiAct [8] explored the correlation between body movement and CSI amplitude to classify activities. Nonetheless, these systems face significant challenges, particularly the limited spatial diversity provided by a small number of antennas.
@@
In typical indoor environments, where multiple rooms are connected to a single Wi-Fi access point (AP), Wi-Fi signals are frequently confronted with low signal-noise ratio (SNR) conditions, alongside interference and multipath effects [10]. Traditional denoising and feature extraction methods, such as principal component analysis (PCA), may not sufficiently address the complexities inherent in Wi-Fi signal data [11]. Our empirical research indicates that PCA exhibits suboptimal performance in activity recognition. This observation underscores the necessity of exploring more sophisticated techniques to improve HAR performance in these environments.
Figure 1: The schematic of the RIS-based human activity recognition system.
In this study, we introduce a novel reconfigurable intelligent surface-assisted human activity recognition (RISAR) method, which is compatible with commercial Wi-Fi devices. RISAR leverages a RIS to capture a wider range of spatially distributed information, thereby enhancing the spatial diversity of Wi-Fi signals [12, 13, 14], as shown in Fig. 1. During the preprocessing phase, we introduce a high-dimensional factor model for extracting activity features from raw CSI measurements in the temporal domain. Additionally, we present a specialized dual-stream spatial-temporal attention network (DS-STAN) architecture tailored to address the nuanced characteristics inherent in sensing data. This architecture is trained to interpret the patterns embedded within CSI data.
@@
  • An innovative RISAR approach is presented for human activity recognition under low SNR conditions. The integration of RIS not only introduces additional links but also enhances spatial diversity, leading to a significant improvement in signal quality and system performance. To the best of our knowledge, this represents the first attempt to leverage RIS technology for human activity recognition using commodity Wi-Fi devices.
@@
  • A high-dimensional factor model approach is proposed for extracting activity features from CSI measurements in the temporal domain. These extracted features are fed into the proposed DS-STAN model, which integrates an attention mechanism to discriminate various human activities. This methodology represents a sophisticated fusion of random matrix theory and deep learning techniques, offering enhanced accuracy in activity recognition.
@@
  • The RISAR system, operational with commercial Wi-Fi devices under various conditions, has undergone extensive evaluation. The empirical results highlight its remarkable accuracy, achieving an average of 97.26% in the Office dataset and 90.83% in the L-shaped hallway dataset. These tests highlight the robustness and adaptability of RISAR across a range of environmental settings.
The remainder of this paper is organized as follows. We start by providing a comprehensive literature review of existing work in Section II. In Section III, we present the system architecture, as well as the preprocessing and recognition algorithms. Section IV illustrates the performance of the proposed approach. Finally, Section V concludes the paper.
II Related Work
In this section, we provide a summary of the relevant literature concerning human activity recognition. There exist three modalities in HAR: vision-based [15, 16, 17, 18], wearable device-based [19, 1, 20, 21], and wireless sensing-based systems [22, 23, 5, 6, 7, 24, 25, 8, 26, 27, 28, 9, 29].
@@ II Related Work
Vision-based methodologies leverage high-resolution cameras to capture images or videos for activity recognition [15]. For instance, Kim et al. [16] introduced a depth video-based HAR system, utilizing skeletal joint features to discern the daily activities of the elderly in indoor settings. Additionally, Franco et al. [17] amalgamated skeleton and RGB data stream analyses to extract features for HAR. These systems are are sensitive to lighting conditions and obstructions, in addition to security and privacy concerns [30].
@@
Wearable device-based HAR approaches utilize technologies such as smartphones [1], IoT devices [19], and sensors [20] for activity detection. Hassan et al. [1] developed a robust HAR system based on smartphone sensor data, capable of distinguishing transitional and non-transitional activities. Bhat et al. [19] employed a low-power IoT device for HAR, pioneering a framework for both online training and inference. Bianchi et al. [20] explored the integration of inertial measurement units (IMUs) with deep learning, where wearable sensors transmit data to a cloud service for activity monitoring. These methods suffer from the drawback of requiring long-term wearing, leading to inconvenience among users [30].
Wireless sensing-based HAR encompasses technologies like radar, radio frequency identification (RFID), and Wi-Fi signals. Radars, capable of detecting human motion without necessitating user-carried devices, offer resilience against variations in lighting conditions due to their broad bandwidth [31]. Zhu et al. [22] proposed a distributed ultra-wideband radar-based system, classifying motions through spatio-temporal patterns of radar Doppler signatures. However, limitations in these systems include portability and hardware costs. RFID systems present a cost-effective alternative, requiring minimal hardware such as chip tags and antennas. Oguntala et al. [23] introduced SmartWall, a system employing passive RFID tags for the recognition of sequential activities. The RFID-based approach offers a cost-effective alternative but encounters robustness issues in complex indoor environments, primarily due to the multipath effect and channel variations [32].
@@
In the evolving landscape of Wi-Fi signal-based technologies, a diverse array of systems have been developed for HAR. Yet, these systems frequently necessitate specialized devices and are often inadequate in non-line-of-sight (NLOS) scenarios. For instance, the WiSee system [5] confirmed the efficacy of Wi-Fi signals in detecting human motion in NLOS and through-the-wall environments, albeit relying on USRP-N210 devices rather than standard commercial Wi-Fi hardware. Yousefi et al. [24] employed long short-term memory (LSTM) networks to encode temporal information during feature learning, setting a new benchmark in HAR using CSI measurements. Similarly, Chen et al. [25] innovatively utilized raw CSI signals in conjunction with an attention-based bi-directional LSTM to ascertain action classes. Feng et al. [26] proposed a three-phase system for multiple human activity recognition, focusing on identifying the start and end points of activities in multi-subject scenarios. Xiao et al. [27] introduced a deep learning-based HAR approach that capitalizes on Wi-Fi signals, transforming segmentation tasks into classification challenges via a convolutional neural network (CNN) algorithm. Guo et al. [28] developed an encoder-decoder framework aimed at mitigating accuracy disparities among individuals in HAR applications. Lu et al. [9] in their CeHAR system, advocated for the amalgamation of CSI amplitude and phase features to enhance action detection efficiency. Lastly, Shalaby et al. [29] presented a deep learning model integrating CNNs, gated recurrent units (GRUs), and LSTMs, designed to extract high-dimensional and time sequence features.
@@
These systems generally do not consider HAR in NLOS environments. This is particularly pertinent in many indoor settings, where multiple rooms often connect to a single wireless signal AP, especially in residential environments. In such scenarios, issues like signal attenuation and multipath fading can significantly impair the accuracy of activity recognition.
III System Architecture
This section focuses on the details of the RISAR. The integration of RIS for HAR not only introduces additional links but also enhances spatial diversity. The architecture of the RISAR system is illustrated in Fig. 2. Following the data flow, the data processing procedure consists of three stages: CSI data collection, data preprocessing, and the human activity classification algorithm.
Figure 2: The architecture of the proposed RISAR system.
III-A Hardware Architecture and Testing Settings
A 1-bit RIS utilized in this study operates at a frequency of 5.8 GHz. It consists of $16\times 10$ reflecting units, with each unit spaced $0.4\lambda$ apart. To simplify the experiments, all devices are positioned at the same height. We utilize two Intel Wi-Fi Link 5300 network interface cards (NICs), serving as both the transmitter (Tx) and receiver (Rx).
To evaluate the effectiveness of RISAR in an extremely low SNR environment, initial tests are conducted within an L-shaped hallway, as depicted in Fig. 3 LABEL:sub@F3-a. The setup involves an Intel 5300 NIC with a horn antenna serving as the Tx, while an Intel 5300 NIC equipped with three omnidirectional antennas acts as the Rx. Positioned at the corner, the RIS functions as a relay to enhance the diversity of the environment. Utilizing the DaS algorithm [33], the RIS is optimized to achieve optimal phase configurations. Its efficiency is validated in our prior study [34], yielding an improvement of at least 10 dBi.
Moreover, to validate the role of the proposed data preprocessing and recognition algorithms, two additional tests are conducted in a meeting room and an office room. Unlike the previous testing scenario in the L-shaped hallway, RISs are not utilized in these setups. This configuration is implemented to isolate the impact of the proposed algorithms. The tests, as illustrated in Fig. 3 LABEL:sub@F3-b, involve data collection in a meeting room covering a $12\times 8.7\,\text{m}^{2}$ area and an office room spanning an $8\times 8.7\,\text{m}^{2}$ space. Both the Tx and Rx are configured with three antennas. The Intel 5300 NIC is set to collect data on the 161-st channel with a 20 MHz channel bandwidth.
(a) L-shaped hallway.
(b) Floor plans of the other two experimental environments.
@@ III-B Data Collection
To capture the instantaneous CSI, which exhibits sensitivity to human activities, this study engaged in the analysis of packets harvested by the Intel Wi-Fi Link 5300 NIC within the temporal domain. Each packet is composed of a matrix characterized by dimensions $N_{\text{Tx}}\times N_{\text{Rx}}\times N_{\text{sc}}$, where $N_{\text{Tx}}$ and $N_{\text{Rx}}$ signify the count of transmitting and receiving antennas, respectively, and $N_{\text{sc}}$ indicates the orthogonal frequency-division multiplexing (OFDM) subcarriers [35]. Specifically, the Intel Wi-Fi Link 5300 NIC outputs CSI data for 30 subcarriers corresponding to each of its three antennas. Consequently, during the tests in the L-shaped hallway, each packet generates a data frame containing $1\times 3\times 30$ CSI measurements. By recording 5000 packets, we obtain a data frame of size $90\times 5000$, representing one activity in the dataset. On the other hand, in the meeting room and office room tests, the data frame consists of $3\times 3\times 30$ CSI measurements. Therefore, each activity is of size $270\times 5000$.
RISAR focuses on recognizing human daily activities. We select 10 typical activities for examination, as delineated in Table I. Participants repeated each activity 100 times to build a comprehensive dataset. Each record is sampled at a rate of 1000 packets per second over a 5-second interval.
@@
L-shaped hallway Meeting room Office